The perimeter of a rectangle is 27 cm,and its area is 35 square cm.Find the length and then width of the rectangle.

Answer by stanbon(75887)

The perimeter of a rectangle is 27 cm,and its area is 35 square cm.Find the length and then width of the rectangle.
-----------
P = 2(L + W)
L*W = 35
---------------------------
L = (P/2)-W
--------
Substitute for "L" and solve for "W":
[(P/2)-W]W = 35
-----------------------------
W^2 - (27/2)W +35 = 0
----------------------------
W = [(27/2) +- sqrt[13.5^2 - 4*35]]/(2
----------------------------------------
W = [13.5 +- 6.5]
Width = 7 cm
---
Since L*W = 35, Length = 5 cm
--------------------------------
Cheers,
Stan H.
--------------